THE ELECTRONIC SPECTRUM OF SiB

\underline{C. R. BRAZIER}, J. I. RUIZ, Department of Chemistry, California State University Long Beach, Long Beach, CA 90840.

An electronic spectrum of the SiB molecule has been observed for the first time. The SiB molecule was produced in a Corona Excited Supersonic Expansion source using a mixture of diborane and silane gases seeded in helium. The emitted light was dispersed with a 1.3m spectrometer and recorded with a CCD camera. The region from 200nm to 800nm was examined and a single electronic system, A^{4}Π − X^{4}Σ^{−} observed with a 0-0 origin at 18432cm^{−1}. Supersonic cooling results in a rotational temperature of 20K and most emission is from the lowest Ω=5/2 component. Observation of the Ω=3/2 component gives the spin-orbit splitting as -23.9 cm^{−1}. The 0-0, 0-1, 0-2, 1-0, 1-2, and 1-3 bands have been analyzed.